Problem 21-5 Statement of cash flows; direct method [LO21-3, 21-8]

Comparative balance sheets for 2018 and 2017 and a statement of income for 2018 are given below for Metagrobolize Industries. Additional information from the accounting records of Metagrobolize also is provided.

METAGROBOLIZE INDUSTRIES
Comparative Balance Sheets
December 31, 2018 and 2017
($ in 000s)
2018 2017
Assets
Cash $ 560 $ 370
Accounts receivable 710 370
Inventory 860 440
Land 700 670
Building 900 900
Less: Accumulated depreciation (200 ) (160)
Equipment 3,400 3,230
Less: Accumulated depreciation (343 ) (300 )
Patent 1,800 2,100
$ 8,387 $ 7,620
Liabilities
Accounts payable $ 960 $ 710
Accrued expenses payable 330 260
Lease liability—land 130 0
Shareholders' Equity
Common stock 3,680 3,500
Paid-in capital—excess of par 800 740
Retained earnings 2,487 2,410
$ 8,387 $ 7,620
METAGROBOLIZE INDUSTRIES
Income Statement
For the Year Ended December 31, 2018
($ in 000s)
Revenues
Sales revenue $ 3,131
Gain on sale of land 80 $ 3,211
Expenses
Cost of goods sold $ 1,160
Depreciation expense—building 40
Depreciation expense—equipment 664
Loss on sale of equipment 25
Amortization of patent 300
Operating expenses 350 2,539
Net income $ 672


Additional information from the accounting records:

  1. Annual payments of $20,000 on the finance lease liability are paid each January 1, beginning in 2018.
  2. During 2018, equipment with a cost of $690,000 (90% depreciated) was sold.
  3. The statement of shareholders' equity reveals reductions of $240,000 and $355,000 for stock dividends and cash dividends, respectively.


Required:
Prepare the statement of cash flows of Metagrobolize for the year ended December 31, 2018. Present cash flows from operating activities by the direct method. (Enter your answers in thousands (i.e., 5,000 should be entered as 5). Amounts to be deducted should be indicated with a minus sign.)
Answer #1

Solution

Metagrobolize Industries

Cash flow statement (Direct Method)

Metagrobolize Industries

Statement of Cash Flows (Direct Method)

For the Year Ended December 31, 2018

Cash from operations:

Cash received from customers

$2,791

Cash paid to suppliers

($1,330)

cash paid to operating expenses

($280)

Cash Provided by Operating Activities

$1,181

Cash from investing activities:

Purchase of equipment

($860)

Sale of equipment

$44

sale of land

$180

Cash Used by Investing Activities

($636)

Cash from Financing Activities:

Cash dividends

($355)

Cash Used by Financing Activities

($355)

Increase in Cash

$190

Beginning Cash Balance

$370

Ending Cash Balance

$560

Computations:

  1. Cash received from customers = sales – increase in accounts receivable

= $3,131 – (710 – 370) = 3,131- 340

Cash received from customers = $2,791

  1. Cash paid to suppliers:

Cash paid to suppliers = cost of goods sold + increase in inventory – increase in accounts payable

Cost of goods sold = $1,160

Increase in inventory = 860 – 440 = $420

Increase in accounts payable = 960 – 710 = $250

Cash paid to suppliers = 1,160 + 420 – 250 = $1,330

  1. Cash paid to operating expenses:

Operating expenses – increase in accrued expenses payable

Operating expenses = $350

Increase in accrued expenses payable = 330 – 260 = $70

Cash paid to operating expenses = $350 - $70 = $280

  1. Land –

Beg. Balance = $670

Land on lease = $130

Total = $800

Less: ending balance = $700

Cost of land sold = $100

Add: gain on sale = $80

Sale proceeds = $180

  1. Equipment –

Beg. Balance = 3,230

Less: cost of equipment sold = 690

Total = $2,540

Less: ending balance = 3,400

Purchase of equipment = $860

Sale proceeds from equipment sales –

Book value of equipment sold = 690 x 10% = $69

Less: loss on sale = $25

Sale proceeds = $44